OpenStudy (ksaimouli):

factor each expression to its simplest form x2-6x

OpenStudy (angela210793):

x(x-6)

OpenStudy (ksaimouli):

x square-6x

OpenStudy (ksaimouli):

question

OpenStudy (ksaimouli):

ok thanks

OpenStudy (angela210793):

x2-6x=x(x-6)

OpenStudy (ksaimouli):

and also x2-3x-10

OpenStudy (angela210793):

x2-3x-10=(x-5)*(x+2)

OpenStudy (ksaimouli):

how to do

OpenStudy (ksaimouli):

how u got